OpenStudy (anonymous):

x^3-8/x^2+2x+4 (x^2-4) i need help please My answers are all partial credit apparently, i dont know what im doing wrong, please help...

OpenStudy (anonymous):

i think x^3-8/x^2+2x+4 divided by (x^2-4) = (x^3-8) / ((x^2+2x+4)(x^2-4)) = ((x-2)(x^2+2x+4)) / ((x^2+2x+4)(x^2-4)) ... factorise x^3-8 = (x-2) / (x^2-4) ... divide numerator and denominator by x^2+2x+4 = (x-2) / ((x+2)(x-2)) ... factorise x^2 - 4 = 1 / (x+2) ... divide numerator and denominator by x-2

OpenStudy (highlyeducated87):

is this the same problem

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

\[\frac{ (x^3-8) }{ ((x^2+2x+4)(x^2-4)) }\] if so you got it